WHAT IS THE MAXIMUM VALUE OF y=-x^2+6x?
This is a quadratic with a=-1 and b=6
The maximum point occurs when x=-b/(2a)
x=-6/(-2)=3
When x=3, y= -3^2+6(3)=9
Maximum at (3,9)
Cheers,
Stan H.

Found 3 solutions by uma, stanbon, jai_kos:
Answer by uma(370)


y = - x^2 + 6x
This represents a parabola opening downward and so the vertex will be the maximum point.
The x co-ordinate of the vertex = -b/2a
= -6/2*(-1)
= -6/-2
= 3
Plugging in x = 3 in the given expression,
y = - 3^2 + 6*3
= -6 + 18
= 12
Thus the maximum value of y = - x^2 + 6x is 18.
Good Luck!!!

Answer by stanbon(75887)

WHAT IS THE MAXIMUM VALUE OF y=-x^2+6x?
This is a quadratic with a=-1 and b=6
The maximum point occurs when x=-b/(2a)
x=-6/(-2)=3
When x=3, y= -3^2+6(3)=9
Maximum at (3,9)
Cheers,
Stan H.

Answer by jai_kos(139)

Given an equation y = -x^2 + 6x
Where a = -1, b = 6
Since the a < 0, we have a maximum value.
x = (-b /2a) = -6 /2 * -1 = 6 /2 = 3
x = 3
Put x =3 in equation(1), we get
y = -(3)^2 + 6 * 3 = -9 + 18 = 9
Therefore the maximum value is given by 9.